what is perpendicular to y = -3x + 7, through (3, 5)?
Answer by mananth(16946)

y = -3x + 7, through (3, 5)?
...
slope = -3
..
the slope of perpendicular = 1/3
..
(3,5) m= 1/3
find b
y = mx +b
5 = 1/3 * 3 +b
5 =1+b
b=4
..
equation : y = x/3 +4
..
...
m.ananth@hotmail.ca